Sol Summary Report – February 18th

Crew 275 Sol Summary Report 18Feb2023

Sol: 6

Summary Title: If the Moon is made of cheese, let Mars be made of brownies!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today, the crew was happy to wake up by the sound of their alarm, rather than by the sound of radio communications with Mission Support. Let’s hope that the temperature will not decrease again and give some very relative rest to the generator (and to Mission Support)! As yesterday’s EVA was postponed to this morning, we already knew the procedures very well and the pre-EVA briefing was relatively short. The three hours of this EVA were used efficiently, as the meteorological mast is now standing proudly right next to Mega-Ares and data acquisition is already ongoing on most of the instruments. In addition, we had time to perform most of the tests on Mega-Ares and launched the acquisition as well, even though some additional tests will be needed during a future EVA (potentially during the deployment of LOAC, that has not been done so far because of the weather conditions). After this successful collaboration between the EVA crew and HabCom, everyone was very happy to share a delicious meal, prepared by the ones who remained at the Hab (we even had brownies!). The other activities continued within the station with absolutely no problem, even though it is always difficult to stick exactly to the hourly schedule meticulously prepared by our Crew Scientist, especially for activities requiring the contribution of several crewmembers.

Look Ahead Plan:

As we are successfully reaching the end of this first week, tomorrow will be a well-deserved rest day for the crew. We will take this opportunity for some team building activities and personal time. This rest time is even more important as week 2 will also be packed with many activities, including the beginning of the photogrammetry EVAs and the end of the deployment of the atmospheric science instruments.

Anomalies in work: None.

Weather: Slight cloud coverage, low -9°C/15°F, high 1°C/34°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site